• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 23:52:28 on April 22, 2008

    Entry number 233831

    Followups:

    keyword=BigBite Camac Controller was reset.

    Paged because of multiple ROC failures when trying to start a CODA run. After several kcoda/startcoda and reboots of ROCs... found that the rate on the E and dE planes was WAY above normal... showing ~5MHz per PMT.

    I found that the Thresholds on the CAMAC discriminators were set to ~10mV... which is a sign that the crate was reset. I set E to 500, and the dE to 280 (as they were before), and re-ran the scripts to set the delay for T3, T4, T7.

    Not sure HOW the CAMAC crate was reset... but I think we're back to normal again. I had the shift workers take a new run... so that I can compare it to previous runs to verify the timing of the triggers.